OpenStudy (anonymous):

Question 3.3. xy = –25. What is the constant of variation for this relationship?

Nnesha (nnesha):

this is equation equation \[\huge\rm y = \frac{ k }{ x }\] where k is constant of varation

Nnesha (nnesha):

solve this for k

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

k=625

Nnesha (nnesha):

hug nope solve that equation for k this one y = k/x

Nnesha (nnesha):

haan. nope solve that equation for k this one y = k/x

OpenStudy (anonymous):

the new equation should be -25=k/-25 right?

Nnesha (nnesha):

hmm nope you don't have to solve anything in this equation there are just 3 variable x y and k multiply both side by x

OpenStudy (anonymous):

ok so yx=k

Nnesha (nnesha):

yes right the equation in the question is same like this one so what is k in that equation ?

Nnesha (nnesha):

k = is constant of variation

OpenStudy (anonymous):

ok so -25

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

ohhh ok thanks so so much!

Nnesha (nnesha):

np :)